19. A swimmer can swim at a speed of 0.70 ms-1 with respect to the water. She wants to cross a river 50 m wide and having a current of 0.50 ms-1. a) In what direction must she swim in order to reach the other bank at a point directly across where she started from? How rapidly will she increase her distance from the near bank? How long will it take her to cross?
b) If instead, she decides to cross in the shortest possible time, how should she swim? Find the time taken to cross?

Grade:11

1 Answers

Aman Bansal
12 years ago

Dear Dennis,

distance = speed * time

downstream
1.2km = speed * 20 minute

speed = 1.2/20 (downstream) = .06 km/min * 1000m/km * 1 min/60sec = 1 m/s downstream

Relative to the stream, 1 m/s - .3 m/s = .7 m/s
therefore
upstream speed = .7 m/s - .3 m/s = .4 m/s

time = distance/speed = 1.2 km / (.4 m/s * km/1000m * 60s/min) = 50 min

total time = 50 + 20 = 70 minutes

Best Of luck

Plz Approve the answer...!!!!

Cracking IIT just got more exciting,It’s not just all about getting assistance from IITians, alongside Target Achievement and Rewards play an important role. ASKIITIANS has it all for you, wherein you get assistance only from IITians for your preparation and winexciting gifts by answering queries in the discussion forums. Reward points 5 + 15 for all those who upload their pic and download the ASKIITIANS Toolbar, just a simple click here to download the toolbar….
So start the brain storming…. become a leader with Elite Expert League ASKIITIANS

Thanks

Aman Bansal

Askiitian Expert